A person running an 800-meter race averages 130 meters per minute for the first 3/4 of the race. The average speed for the remainder of the race is 145 meters per minute. What is the person's average speed for the entire 800 meters rounded to the nearest whole number?
Answer by mananth(16949)

A person running an 800-meter race averages 130 meters per minute for the first 3/4 of the race. The average speed for the remainder of the race is 145 meters per minute. What is the person's average speed for the entire 800 meters rounded to the nearest whole number?
..
3/4 *800= 600 meters
Avg speed = 130 mtr/minute
time taken = 600 /130 minutes
..
next 200 meters
speed = 145m/minute
time = 200/145 minutes
..
Total time taken = 600/130 + 200/145
5.994 minures = 6 minutes
..
total distance = 800 meters
total time = 6 minutes
average speed = 800/6 = 133.33 meters/minute